STRUCTURE AND INVERSION MOTIONS OF THE WEAKLY BOUND CH_{2}F_{2} ⋯CO_{2} DIMER

\underline{REBECCA A. PEEBLES}, AMELIA J. THOMAS, MICHAL M. SERAFIN and SEAN A. PEEBLES, Department of Chemistry, Eastern Illinois University, 600 Lincoln Ave., Charleston, IL 61920.

The rotational spectrum of the CH_{2}F_{2} ⋯CO_{2} dimer has been measured using chirped-pulse and resonant cavity Fouriertransform microwave (FTMW) spectroscopy, with the broadband spectrum playing an essential role in allowing identification of the tunneling splittings. The observed a-type transitions were doubled by a few megahertz, while c-type transitions were split by around 200 MHz, suggesting a tunneling motion that inverts the μ_{c} dipole moment component. A Pickett-type coupled Hamiltonian has been used to analyze the spectra, giving an energy difference (ΔE) between the tunneling states for the normal isotopologue of 115.140(2) MHz, and rotational constants of A_{0} = 5567.8604(27) MHz, B_{0} = 1832.9676(5) MHz, C_{0} = 1828.6132(5) MHz, A_{1} = 5567.8540(26) MHz, B_{1} = 1831.7711(4) MHz, C_{1} = 1828.6106(4) MHz. It was also necessary to include a Coriolis coupling term (G_{b} = 7.740(6) MHz) as well as fourth and sixth order centrifugal distortion constants to obtain a satisfactory spectroscopic fit. The rotational constants and planar moments are consistent with a C_{s} symmetry structure in which the C_{2} axis of CH_{2}F_{2} makes an angle of roughly 23^{∘} with the axis of the CO_{2}, with the fluorine atoms of CH_{2}F_{2} straddling the CO_{2} carbon atom. The spectra of five additional isotopologues were analyzed, providing detailed structural information, and all except the mixed C^{18}O^{16}O species showed inversion splittings similar to the normal species. The observation of unsplit spectra for two distinct C^{18}O^{16}O isotopologues confirms that the internal motion involves movement of the CH_{2}F_{2} subunit between the two ends of the CO_{2} molecule.